OpenStudy (anonymous):

A jet is flying with a uniform motion of 9.3 x 10^2 ms/ (S), magnitude approx. Mach 2.7. At time zero it passes a a mountain top. Construct a table

OpenStudy (anonymous):

where are you needing help at? and is "9.3 x 10^2 ms/ (S)" meaning milliseconds per Second?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

How do I construct the table to show the plane's position to the mountain top at the end of each second for a 12 s period?

OpenStudy (anonymous):

what is the "(S)"?

OpenStudy (anonymous):

I dont exactly get what they're trying to ask... :( and S= seconds

OpenStudy (anonymous):

I think I get it, but its taking me a second to gather my thoughts to make sure I understand what they quesiton is really asking ^_^

OpenStudy (anonymous):

Its alright :)

OpenStudy (anonymous):

Are you sure (S) isn't south and ms isn't actually m/s ?

OpenStudy (anonymous):

Omg, yes you're right. I'm so sorry!!! I typed it out to quickly!!!

OpenStudy (anonymous):

mach 1 = speed of sound = 295.1 m/s so mach 2.7 = (295.1) * 2.7 = 796.8 m/s and you're told 930 m/s? I'm not sure what these two speeds mean... or how thy relate...

OpenStudy (anonymous):

Yes thats what it is said in the book.... 9.7 x 10^2 m/s (S) :(

OpenStudy (anonymous):

Does it give another direction for the mach 2.7 speed?

OpenStudy (anonymous):

boy... i'm not sure which speed to use.. maybe make a table with both speeds? or does one speed mean south and the other speed means east or west?

OpenStudy (anonymous):

No it doesn't give another direction... :(

OpenStudy (anonymous):

How did you get the 295.1 m/s by the way?

OpenStudy (anonymous):

mach 1 means speed of sound which equals 295.1 m/s ^_^

OpenStudy (anonymous):

Mach 1 is the speed of sound (Aprox. 295.1 m/s ). To get Mach Whatevernumberyouwant, you just multiply Mach 1 times whatevernumberyouwant. In this case Mach 2.7 = 295.1*2.5 m/s.

OpenStudy (anonymous):

Did you research that? Sorry :$

OpenStudy (anonymous):

Is this suppose to be known on that top of my head or was it researched? Sorry, I'm just trying to figure it out :$ D:

OpenStudy (anonymous):

Ehmmm, most people know that the speed of sound is approximately 300 m/s. It's a sort of constant (It really isn't, but it doesn't change that much). It's just like memorizing that pi is 3.1416.

OpenStudy (anonymous):

Ohh alright, thank you :) So I would need to construct two tables then?

OpenStudy (anonymous):

I dont exactly get how this would look like :( Im really confused...

OpenStudy (anonymous):

I think constructing atleast the one table using the speed of 930 m/s is best rout since the question is unclear..

OpenStudy (anonymous):

"A military jet is flying with uniform motion at 9.3×10^2 m/s [S], the magnitude of which is approximately Mach 2.7. At time zero, it passes a mountain top, which is used as the reference point for this question. " I'd guess you just have to tabulate for 930 m/s. Which, considering that the speed is constant, you just have to multiply 930 times every integer between 0 a 12.

OpenStudy (anonymous):

(ALSO, fun fact! (930 m/s)/2.7 is 344 m/s, which is the speed of sound at sea level. So the jet would need to be traveling on land so that it'd actually be traveling at Mach 2.7)

OpenStudy (anonymous):

I researched it back, and here is how the two number relate at an altitude of sea level, Mach 1 = 340 m/s, if we take the 930 / 340 = 2.73 so (Lessis beat me to it ^_^)

OpenStudy (anonymous):

@Lessis you got this, i'm gonna help another

OpenStudy (anonymous):

So would the table be 0 = 0, 1 = 930, 2= 1860...?

OpenStudy (anonymous):

Considering it's speed constant, yes. \[V = \frac{ d }{ t }\] So\[d = Vt\] So the distance at t = 1 second would be V*1. And at 2 V*2 and so on.

OpenStudy (anonymous):

Do I multiply 930 m/s by 0,1,2,3 so on?

OpenStudy (anonymous):

Up to 12 m/s right?

OpenStudy (anonymous):

Yes, and no. t=0, t= 1, t= 2, t=3, t=4, ... , t=12 are seconds, not speed. And the result you get is a distance, in meters. (d = Vt)

OpenStudy (anonymous):

What you're doing is correct, just check your units so you get what is asked and don't screw up. (Distance in meters, speed in meters/second, time in seconds, etc)

OpenStudy (anonymous):

Ohhhhhh... Okay :))

OpenStudy (anonymous):

Thank you so much! :)

OpenStudy (anonymous):

You're welcome. ;D

OpenStudy (anonymous):

I'm sorry it took awhile... I'm a slow learner :(